Demand

Complete the table bellow, if demand fuction is Q = 250 – 25P 



Questions:

a. Describe the relationship between demand, TR, MR and price elasticity!

b. At what price the demand is inelastic, elastic, & unitary?

c. When demand is elastic and inelastic, should the price be lowered or should it be increased? Prove it with the numbers from the table.
Answer #1

Price(harga) Quantity (Q) TR=P*Q MR=ATR/AQ Ed 10 0 8 50 400 8 -9.00 75 525 5 -3.00 6 100 600 -1.86 5 125 625 1 -1.22 4 150 60no + (04-0.) 2 to aindd 9) Ed = % A in P (-1)/IP 2 P($) esi Je=1 ecl UMR dd. 9 TRIB) т TR curve > Q b) when ledl> 1 then theine c) when the elastic demand is elastic ť in TR lead to then tin Price 0 dd. is inelastic wherean it lead to A in TR ↑ in P